i need help on two problems if its possible!...
5-(square root 3x)=1
4x=1(all in sqaure root)=3
thanx
Trish

Answer by mukhopadhyay(490)

5-(square root 3x)=1
=> (square root 3x)=5-1
=> (square root 3x)=4
=> 3x=16
=> x=16/3;
Please check your other problem if copied correctly